1.Suppose a consumer consuming two commodities X and Y has the following utility function TU= X0.5 Y0.5. If price of good X and Y are 2 and 3 respectively and income constraint is Birr 60.


A. Formulate the budget equation


B. Find the MRSXY at optimum.


C. Find the quantities of good X and Y which will maximize utility

1
Expert's answer
2022-02-04T06:37:02-0500

"P_x= 2"

"P_y= 3"

I= 60

a) Budget Eqn

2X+ 3Y= 60


b) MRS"_{xy}= \frac{Mu_x}{Mu_y}"

"Mu_x= 0.5X^{-0.5}Y^{0.5}"


"Mu_y= 0.5 X^{0.5}Y^{-0.5}"


MRS"_{xy}= \frac{ 0.5X^{-0.5}Y^{0.5}}{0.5 X^{0.5}Y^{-0.5}}"


"= \frac{ 0.5Y^{0.5}Y^{0.5}}{0.5 X^{0.5}X^{0.5}}= \frac{Y}{X}"

c) Quantities that maximize utility


"\frac{Mu_x}{Mu_y}=\frac {P_x}{P_y}"


"\frac{Y}{X}=\frac {2}{3}"

3Y= 2X

"X= \frac{3}{2}Y"


"Y= \frac{2}{3}X"

Plug into the budget equation

2("\frac{3}{2}Y)+ 3Y= 60"


6Y= 60

Y"^*= 10"


"3(\frac{2}{3}X)+ 2X= 60"

4X= 60

X"^*= 15"